Displaying report 1-1 of 1.
Reports until 17:51, Tuesday 30 July 2024
Atoms General
Briana Chen - posted 17:51, Tuesday 30 July 2024 (11791)
Locking!!!

[Briana, Ian, Torrey]

Before exciting stuff: This morning, I noticed that photodetector saturation was occurring at 3.3 mW. All settings were the same as yesterday (Input 1 channel at 50 Ohms, 40 Vpp). The way that I returned to yesterday's voltage level was to switch to 1 MOhm and then switch back to 50 Ohm. For some reason, this gets you back to the higher voltage level. You also have to be at a high-enough power (>5 mW I want to say? Certainly 8 mW) for the lack of change between 1 MOhm and 50 Ohms to occur. I have no idea why- the change from 1 MOhm to 50 Ohm should change the voltage reading by a factor of 1/2. Either way, this process returned it to the voltage level from yesterday (~6 V). 

Retook some dip depth measurements to check repeatability- it seems repeatable within the error bars (power error bars are too small to be seen): dips.png (red and purple lines were taken today). Even at a different current, the saturation tends to occur around the same place. This is plotting absolute dip depth, which will help us determine at what power saturation occurs. Data is here: 7_26.xlsx

Shoutout to Torrey for figuring out that we need a high-bandwidth photodetector to detect the higher frequency signals, so we needed a PDA10A2 detector (150 MHz) instead of the PDA36A2 detector (12 MHz). We exchanged our 36A2 with Torrey's 10A2. The 10A2 has an active area of about 0.8 mm^2 (1 mm diameter ish), so we may need to add a lens to focus it down. We also added two ND filters (a 2 OD and 0.3 OD filter) so that only ~0.166 mW hits the detector. Both filters should be placed right before the photodetector so the light going through the vapor cell is still a high power. Will call Thorlabs to ask what the damage threshold is, but Torrey remembers it being around 1 mW, which is why we added the ND filters. 

The confusion with the attenuation of -40 dB automatically by the multi-instrument mode means that the signal in single instrument mode should be multiplied by 0.01 (-40 dB = 10 log (P2/P1) = 10 log (V2^2 / V1^2) so V2 = 0.01 * V1). This is approximately the case: mimvssim.png. There doesn't seem to be a way to change this, it may just be the way the MIM operates. It is probably fine too since we have a low power hitting the photodetector now anyways. 

Locking!:

We locked the laser to the main absorption dip!! See pictures of the error signal (errorsignal.png) and the result when locked (locked.png). It seems like the controller should operate at lower frequencies (15 Hz) to best lock, which Torrey pointed out is different from the kHz frequencies he is using for the cavity locks and means we could consider using the slow controller. Torrey definitely did not lock the laser or help at all. The next step is to better understand the locking process before Ian deletes it all. Also, should tune controllers/modulation frequency/phase to make sure the error signal is maximized. Setup is same as in the picture last post except now with a different photodetector and some ND filters. 

Images attached to this report
Non-image files attached to this report
Displaying report 1-1 of 1.